OpenStudy (anonymous):

Given b = 4 and h = 1, what is the equation of the graph if the parent function is y=√(x)

OpenStudy (anonymous):

hmmm....

OpenStudy (anonymous):

what a bizarre and idiosyncratic question this is not any standard math i have seen before

OpenStudy (anonymous):

what is \(b\)? what is \(h\)? i mean "what do they stand for?"

OpenStudy (anonymous):

So assuming b as base, and h as height.

OpenStudy (anonymous):

is it something like \(f(bx+h)\)? you need to look back and see how they are used

OpenStudy (anonymous):

i really doubt it is base and height

OpenStudy (anonymous):

You need more information for this problem.

OpenStudy (anonymous):

yes but you need to go to the book or whatever source you are using and see exactly what it says, otherwise we are just guessing

OpenStudy (anonymous):

all what i can get is y=asqrt4(x-1) +k

OpenStudy (anonymous):

So it would probably be A

OpenStudy (anonymous):

nothing like made up math is there? where does this question come from? some on line source? a book? a teacher?

OpenStudy (anonymous):

if this was made up by a teacher, frankly she should not be teaching this would be like me asking you if \(f(x)=x^2\) , \(a=4,b=3\) what do you get? the question makes literally no sense out of context, and there is no standard context i know of that uses \(b\) and \(h\) as specific variables

OpenStudy (anonymous):

So was the answer correct?

OpenStudy (anonymous):

this one does make sense \[g(f(x))=g(2x-1)=2x-1+2=2x+1\]

OpenStudy (anonymous):

ok, well let me know if its right or wrong, I would like to know.

OpenStudy (anonymous):

second one i answered above, hope steps are clear